Let X be the number of material anomalies occurring in a particular region of an aircraft gas-turbine disk. A researcher proposes a Poisson distribution for X. Suppose that ? = 6 The Poisson probability mass function is: P(x-fr 0,1,2.. Use the pmf to calculate probabilities. Verify these values in R using dpois(x,lambda) Compute the following probabilities: (Round your answers to three decimal places.) (a) P(X-3)- (c) P(X< 3) (d) PX 3)-
Answer #1

e_λλκ e-6(6)3 (0.002 479) × (216) P(X = 3) =--31 Use excel function POISSON.DIST( X, mean, cumulative) to calculate Px(3): PO0.002 479 0! 0! 0.014 873 = 1! P(X = 2) = 6(6)2 Answer 0.002 479 + 0.014 873 + 0.044 618 0.061 969 Use excel function POISSONUse excel function POISSON.DIST( X, mean, cumulative) to calculat<e P(X > 3)-1-P(X < 3) = 1-Ex (3): 1-POISSON.DIST( 3, 6, TRU